# Neural Transparency: Mechanistic Interpretability Interfaces for Anticipating Model Behaviors for Personalized AI

### v1

@@ Abstract.
Millions of users now design personalized LLM-based chatbots that shape their daily interactions, yet they can only loosely anticipate how their design choices will manifest as behaviors in deployment. This opacity is consequential: seemingly innocuous prompts can trigger excessive sycophancy, toxicity, or inconsistency, degrading utility and raising safety concerns. To address this issue, we introduce an interface that enables neural transparency by exposing language model internals during chatbot design. Our approach extracts behavioral trait vectors (empathy, toxicity, sycophancy, etc.) by computing differences in neural activations between contrastive system prompts that elicit opposing behaviors. We predict chatbot behaviors by projecting the system prompt’s final token activations onto these trait vectors, normalizing for cross-trait comparability, and visualizing results via an interactive sunburst diagram. To evaluate this approach, we conducted an online user study using Prolific to compare our neural transparency interface against a baseline chatbot interface without any form of transparency. Our analyses suggest that users systematically miscalibrated AI behavior: participants misjudged trait activations for eleven of fifteen analyzable traits, motivating the need for transparency tools in everyday human-AI interaction. While our interface did not change design iteration patterns, it significantly increased user trust and was enthusiastically received. Qualitative analysis indicated that users’ had nuanced experiences with the visualization that may enrich future work designing neurally transparent interfaces. This work offers a path for how mechanistic interpretability can be operationalized for non-technical users, establishing a foundation for safer, more aligned human-AI interactions.
@@ 1. Introduction
Human-AI interaction has become increasingly personalized and ubiquitous with the rise of customizable AI companions powered by large language models (LLMs)(Pataranutaporn et al., 2021; Siemon et al., 2022; Ta et al., 2020; De Freitas et al., 2024; Patel, 2024; Pataranutaporn et al., 2025; Mahari and Pataranutaporn, 2025; Fang et al., 2025; Liu et al., 2024). Platforms like Character.AI have attracted over 20 million monthly active users worldwide, who have collectively created 18 million unique chatbots (Kirk et al., 2025). This unprecedented scale of AI companion creation reflects a fundamental shift: users no longer passively interact with pre-configured assistants but actively design AI personalities tailored to their specific needs, preferences, and relationships. These custom chatbots have become deeply integrated into users’ lives, serving as confidants, creative collaborators, study partners, and emotional support systems (Mahari and Pataranutaporn, 2025; De Freitas et al., 2024; Liu et al., 2024). The intimacy of these relationships, with users spending hours daily conversing with their created companions, means that chatbot behaviors carry significant weight in shaping users’ emotional well-being, decision-making, and worldviews (De Freitas et al., 2024; Fang et al., 2025; Phang et al., 2025; Costello et al., 2024).
@@
However, this creative freedom comes with substantial risks. Even minor modifications to system prompts — the foundational instructions that configure a model’s behavior and persona before any user interaction begins — can trigger unintended and problematic behaviors. For instance, a small addition to ChatGPT’s system prompt in 2025 resulted in such sycophantic responses that widespread user complaints forced OpenAI to swiftly roll back the change (OpenAI, 2025). As users craft their own system prompts to shape their chatbot’s personality, they often trigger model behaviors they neither anticipated nor intended (Zhang et al., 2025). A seemingly innocuous instruction to ”be supportive” might inadvertently produce extreme sycophancy (Sharma et al., 2023; Malmqvist, 2025; Ibrahim et al., 2025), where the chatbot never challenges harmful ideas. A prompt designed to create an ”edgy” personality might cross the line into promoting toxicity or violence (Zhang et al., 2025). These emergent behaviors are particularly concerning given recent reports of AI-related psychological harm, including cases of ”AI psychosis” (Yeung et al., 2025; Morrin et al., 2025; Fieldhouse, 2023; Dohnány et al., 2025; Østergaard, 2023) (where vulnerable users lose touch with reality through maladaptive chatbot interactions) and tragic incidents of teenagers taking their own lives following intense relationships with AI companions (Mahari and Pataranutaporn, 2025). The stakes are especially high for adolescent users, who represent a significant portion of the user base and may lack the critical distance to recognize problematic interaction patterns.
@@
The core challenge is the behavioral predictability of LLMs: users currently have no way to anticipate how their design choices will manifest in actual chatbot behavior until they deploy and test their creation. Even then, problematic behaviors may only emerge in specific conversational contexts that users never explore during testing. This ”black box” creation process forces users into reactive mode, discovering issues after they’ve already occurred rather than proactively designing around them (Chao et al., 2025). The problem is compounded by the complexity of modern LLMs, where subtle changes in prompting can produce dramatic shifts in personality (Choi et al., 2024; Frisch and Giulianelli, 2024; Tosato et al., 2025), and where the same prompt can elicit different behavioral profiles across model versions or architectures (Li et al., 2024).
@@
Mechanistic interpretability (MI) offers a promising path forward. Unlike traditional explainable AI approaches that focus on post-hoc rationalizations of model outputs, mechanistic interpretability investigates the causal structure of neural networks by analyzing activation patterns within the model itself (Conmy et al., 2023; Sharkey et al., 2025; Nanda et al., 2023a; Cywiński et al., 2025). Recent work has shown that LLM representations encode rich semantic information about persona, sentiment, and behavioral tendencies in interpretable linear spaces (Cywiński et al., 2025; Chen et al., 2025). By examining how input tokens influence internal activations, researchers have uncovered directions in the activation space that correspond to specific traits, and have shown that these can be manipulated to control model behavior (Chen et al., 2025). Furthermore, MI techniques have revealed how models internally represent their perception of the user, suggesting that chatbots maintain implicit models of who they’re talking to that shape their responses (Viégas and Wattenberg, 2023; Chen et al., 2024).
@@
Despite these theoretical advances, mechanistic interpretability has remained largely confined to AI research communities (Sharkey et al., 2025). Such techniques are mathematically sophisticated and require specialized expertise to apply, and so have not been translated into practical tools that end-users can leverage. We introduce the concept of neural transparency, an interface design approach that translates neural-level model behaviors into interpretable, actionable feedback for non-technical users. Unlike post-hoc explainability approaches that rationalize outputs after the fact, neural transparency exposes predictive insights about behavior before deployment, enabling users to make informed design decisions based on internal representations. This paper bridges the gap between mechanistic interpretability and human-AI interaction, highlighting how neural-level insights can be operationalized to support more informed chatbot creation. To our knowledge, this work is one of the first studies to bring MI techniques directly into user-facing AI tools.
@@
Our Approach We present a novel neural transparency interface for LLM-based chatbot creation that analyzes neural activation patterns to provide real-time predictions of personality traits resulting from custom system prompts. We use an LLM to generate contrastive behavioral examples to create persona vectors (linear representations of binary behavioral traits within the model’s neural activation space). As users craft and refine their system prompts, our interface computes persona scores across behavioral dimensions spanning both desirable traits (empathy, humor, sociality, encouraging, formality) and concerning unsafe behaviors (sycophancy, toxicity, hallucination). The interface presents these predictions through intuitive visualizations, allowing users to see how their design choices might manifest across different interaction contexts before actually talking to their chatbot. Critically, users can iterate on their system prompts and immediately observe how changes affect predicted behaviors, enabling a exploratory and mechanisitically informed design process. Rather than discovering the resulting personality and potential problems through trial and error after deployment, users can proactively identify and mitigate risks during the creation phase.
@@
We conducted a controlled study to evaluate how mechanistic interpretability-based feedback influences users’ comprehension of chatbot behaviors and their prompt refinement strategies. Our study examines:
@@
  • whether neural transparency feedback improves user comprehension of their personalized AI, shapes their perception of the model, and helps them achieve their desired system prompt;
@@
  • how accurately users can anticipate their AI’s behavior compared to ground truth model activations;
@@
    Empirical evidence demonstrating how mechanistic interpretability-based feedback improves user trust in chatbots while also being perceived as useful and interactive.
@@ 2.1. Personality and Safety of Personalized AI
As AI systems become increasingly personalized, ensuring safe interaction while preserving user creative control presents a fundamental tension. Some of these current safety techniques include post-training methods that incorporate human and AI feedback on model outputs (Christiano et al., 2017; Bai et al., 2022; Rafailov et al., 2023), evaluation against safety benchmarks measuring behaviors such as truthfulness (Lin et al., 2021) and toxicity (Hartvigsen et al., 2022), and adversarial testing through red-teaming exercises (Ganguli et al., 2022).
However, current safety mechanisms may be insufficient as the non-deterministic nature of LLMs can lead to user-modified prompts producing unpredictable behavior (Wang et al., 2023; Kim et al., 2024; Ibrahim et al., 2025; Fitz et al., 2025). Recent work has shown that training language models to exhibit warm and empathetic personas, traits users commonly desire, undermines the AI’s reliability and increases error rates (Ibrahim et al., 2025). This finding reveals that even seemingly benign personality traits can create systematic safety risks that standard evaluation practices may fail to detect. The personality given to an LLM using the system prompt has additionally been shown to influence model performance on safety benchmarks (Fitz et al., 2025; Zheng et al., 2024), raising potential safety concerns. Current characterization of these chatbot personalities has been done using established frameworks such as the big five personality trait model (Jiang et al., 2022, 2023), but such traits may be less relevant to user preferences. However, existing methods to assess model personality operate at inference time by analyzing model responses, creating two significant limitations: (1) prevention of rapid iteration during system prompt design and (2) requirement of substantial computational resources in each round of inference. Other work has explored user agency with model personality by matching users with pre-defined LLM personas for support (Tu et al., 2023) and incorporating role-playing personas to enhance zero-shot reasoning (Kim et al., 2024). Despite these advances, limited research exists on enabling users to design custom personas tailored to their specific needs or on developing interfaces that facilitate this design process.
@@ 2.2. Mechanistic Interpretability
There is increasing evidence that LLMs represent features as linear directions in the representational space created by its activations (Elhage et al., 2022; Nanda et al., 2023b; Zou et al., 2023). This phenomenon arises from polysemanticity, whereby LLMs encode more features than available neurons, necessitating individual neurons to represent multiple features through linear combinations of their activation values (Elhage et al., 2022). Various concepts and behaviors, such as refusal (Arditi et al., 2024), sentiment (Tigges et al., 2023), truth (Marks and Tegmark, 2023), political beliefs (Kim et al., 2025), and spatial and temporal relationships (Gurnee and Tegmark, 2023) appear to be encoded in this way. Personality traits can also be encoded as linear directions through persona vectors, which can be found by using difference-in-means between contrastive model responses(Chen et al., 2025).
Linear probes (Alain and Bengio, 2016), which use linear classifiers applied to activations of the model, can similarly identify and manipulate these linearly represented features. Previous work has used linear probes to build human-AI interfaces that measure an LLM’s internal representations of user demographics along these linear feature dimensions (Chen et al., 2024). The interface allowed users to both directly view and manipulate these internal representations and was found to improve transparency and user experience. However, linear probes require extensive data collection and classifier training, whereas persona vectors can be computed from smaller datasets and without requiring model training.
While these methods enable model interpretability, translating their insights into actionable user feedback requires effective visualization. Prior work in using MI techniques to visualize the internal representations and mechanisms of AI includes Neuronpedia (Lin, 2023), an open source repository and platform for researchers and others with a technical background. Among the most notable features is one that allows users to explore the features present in the activations of LLMs using sparse autoencoders (SAEs) (Cunningham et al., 2023). SAEs are also applied in other tools that allow for activation steering (Turner et al., 2023), where the user can directly manipulate the values of the activations to change behavior. Another is circuit tracing (Ameisen et al., 2025), which allows users to view and analyze the connections between features to reveal the LLM’s internal reasoning process. These tools simplify the application of different methodologies in MI to allow researchers to explore the complex internal representations of LLMs with a user interface. While these tools represent initial attempts to leverage MI for making AI mechanisms more accessible, a critical gap remains: translating these insights for non-technical users to enhance their understanding and control over AI systems.
@@ 3. Methodology
Our neural transparency approach translates insights from mechanistic interpretability into a practical tool for chatbot creation. The core design challenge was making neural activation patterns, typically only analyzed by AI researchers, interpretable and actionable for non-technical users. We developed a web-based interface that uses persona vectors extracted from model activations to provide real-time predictions about chatbot behavioral traits before users deploy and interact with their creation.
We chose chatbot companion creation as our application domain as it represents a high-stakes context where behavioral prediction failures can have serious consequences (Mahari and Pataranutaporn, 2025; Fang et al., 2025; Zhang et al., 2025; Liu et al., 2024; De Freitas et al., 2024; Kirk et al., 2025), yet users typically receive no transparency about how their design choices translate into AI behavior. We focused specifically on emotional support chatbots, a use case that balances ecological validity (millions use AI companions for emotional support) (Kirk et al., 2025) with ethical constraints (avoiding explicitly adversarial tasks in a controlled study). This task requires users to balance multiple competing objectives: warmth and empathy without excessive sycophancy, honesty without coldness—making it an ideal testbed for evaluating whether neural transparency helps users navigate complex tradeoffs.
@@ 3.1.1. Generating Prompts for Trait Extraction
The first step to create persona vectors was generating contrastive system prompts and situation questions that extract the selected behavior from our target LLM, Llama-3.2-3B-Instruct. We chose eight personality dimensions (empathy, sociality, encouraging, funniness, formality, sycophancy, hallucination, and toxicity), each with opposing trait pairs for a total of 16 traits. The first five were selected for relevance to personality traits a user might desire in an AI chatbot. The final three (sycophancy, hallucination, and toxicity) were selected to inform the user about the safety implications of their system prompt.
We generate five contrastive pairs of system prompts using Claude Haiku 3.5. Each pair contained a positive system prompt that instructs the model to exhibit the selected trait and a negative system prompt that instructs the model to exhibit the opposite of the selected trait. These were to be used to find the direction in the activation space that control the expression of different levels of the target trait. An example for empathy is the following:

### v2

@@ Abstract.
Millions of users now design personalized LLM-based chatbots that shape their daily interactions, yet they can only roughly anticipate how their design choices will manifest as behaviors in deployment. This opacity is consequential: seemingly innocuous prompts can trigger excessive sycophancy, toxicity, or other undesirable traits, degrading utility and raising safety concerns. To address this issue, we introduce an interface that enables neural transparency by exposing language model internals during chatbot design. Our approach extracts behavioral trait vectors (empathy, toxicity, sycophancy, etc.) by computing differences in neural activations between contrastive system prompts that elicit opposing behaviors. We predict chatbot behaviors by projecting the system prompt’s final token activations onto these trait vectors, normalizing for cross-trait comparability, and visualizing results via an interactive sunburst diagram. To evaluate this approach, we conducted an online user study using Prolific to compare our neural transparency interface against a baseline chatbot interface without any form of transparency. Our analyses suggest that users systematically miscalibrated AI behavior: participants misjudged trait activations for eleven of fifteen analyzable traits, motivating the need for transparency tools in everyday human-AI interaction. While our interface did not change design iteration patterns, it significantly increased user trust and was enthusiastically received. Qualitative analysis revealed nuanced user experiences with the visualization, suggesting interface and interaction improvements for future work. This work offers a path for how mechanistic interpretability can be operationalized for non-technical users, establishing a foundation for safer, more aligned human-AI interactions.
@@ 1. Introduction
Human-AI interaction has become increasingly personalized and ubiquitous with the rise of customizable AI companions powered by large language models (LLMs) (pataranutaporn2021ai; siemon2022we; ta2020user; de2025ai; patel2024chatbot; pataranutaporn2025my; mahari2025addictive; fang2025ai; liu2024chatbot). Platforms like Character.AI have reported over 20 million monthly active users worldwide (characterai2025warpstream), with users creating more than 2.7 million personalized chatbots in the process (characterai2023funding). This unprecedented scale of AI companion creation reflects a fundamental shift: users no longer just passively interact with pre-configured assistants but actively design AI personas tailored to their specific needs, preferences, and relationships. These custom chatbots have become deeply integrated into users’ lives, serving as emotional support, creative collaborators, personal assistants, and even romantic partners (de2025ai; liu2024chatbot). The intimacy of these relationships, with users spending hours daily conversing with their created companions, means that chatbot behaviors carry significant weight in shaping users’ emotional well-being, decision-making, and worldviews (de2025ai; fang2025ai; phang2025investigating; costello2024durably).
@@
However, this creative freedom comes with substantial risks. Even minor modifications to system prompts—the core instructions that configure a model’s behavior and persona before any user interaction begins—can trigger unintended and problematic behaviors. As users craft system prompts to shape their chatbot’s persona, they can trigger model behaviors they neither anticipated nor intended (zamfirescu2023johnny). A seemingly innocuous prompt to ’be supportive’ might inadvertently produce sycophancy (sharma2023towards; malmqvist2025sycophancy; ibrahim2025training), where the chatbot never challenges harmful ideas. Similarly, a prompt to role-play an ‘edgy’ persona might cross the line into promoting toxicity or violence (zhang2025dark). These emergent, unintended behaviors are particularly concerning given recent reports of AI-related psychological harm, including cases of AI-induced psychosis (morrin2025delusions; fieldhouse2023can; dohnany2025technological). In such cases, vulnerable users may lose touch with reality through maladaptive role-play-based interactions, leading to tragic incidents such as teenagers taking their own lives following intense relationships with AI companions (mahari2025addictive). The stakes are especially high for adolescent users, who represent a significant portion of the user base and may lack the critical distance to recognize problematic interaction patterns.
@@
The core challenge is the behavioral predictability of LLMs: users currently lack principled ways to anticipate how their persona design choices will manifest in actual chatbot behavior until they deploy and test their creation. Even then, problematic behaviors may only emerge in specific conversational contexts that users can only discover after deployment, discovering issues after they’ve already occurred rather than proactively designing around them (chao2025jailbreaking). The problem is compounded by the complexity of modern LLMs, where interaction and changes in prompting can produce dramatic shifts in personality (frisch2024llm; tosato2025persistent), and where the same prompt can elicit different behavioral profiles across model versions or architectures (li2024measuring).
@@
Mechanistic interpretability (MI) offers a promising path forward. Unlike traditional explainable AI approaches that focus on post-hoc rationalizations of model outputs, mechanistic interpretability investigates the causal structure and patterns within neural networks (bereska2024mechanistic; sharkey2025open). Recent work has shown that LLM representations encode rich semantic information about persona, sentiment, and behavioral tendencies in interpretable linear spaces (tigges2023linear; chen2025persona). By examining how input tokens influence internal activations, researchers have uncovered directions in the activation space that correspond to specific traits, and have shown that they can be manipulated to control model behavior (chen2025persona). Furthermore, MI techniques have revealed how models internally represent their perception of the user, suggesting that chatbots maintain implicit models of who they’re talking to that shape their responses (chen2024designing).
@@
Despite these theoretical advances, mechanistic interpretability has remained largely confined to AI research communities (sharkey2025open). Such techniques are mathematically sophisticated and require specialized expertise to apply, and so have not been translated into practical tools that end-users can leverage. We introduce the concept of neural transparency, an interface design approach that translates neural-level model behaviors into interpretable, actionable feedback for non-technical users. Unlike post-hoc explainability approaches that rationalize outputs after the fact, neural transparency exposes predictive insights about behavior before deployment, enabling users to make informed design decisions based on internal representations. This paper bridges the gap between mechanistic interpretability and human-AI interaction, highlighting how neural-level insights can be operationalized to support more informed chatbot creation. To our knowledge, this work is one of the first studies to bring MI techniques directly into user-facing AI tools.
Our Approach
@@
We present a novel neural transparency interface for LLM-based chatbot creation that analyzes neural activation patterns to provide predictions of personality traits resulting from custom system prompts. We use an LLM to generate contrastive behavioral examples to create persona vectors (linear representations of binary behavioral traits within the model’s neural activation space). As users craft and refine their system prompts, our interface computes persona scores across behavioral dimensions spanning both desirable traits (empathy, humor, sociality, encouraging, formality) and unsafe behaviors (sycophancy, toxicity, hallucination). The interface presents these predictions through intuitive visualizations, allowing users to see how their design choices might manifest across different interaction contexts before actually talking to their chatbot. Critically, users can iterate on their system prompts and immediately observe how changes affect predicted behaviors, enabling an exploratory and mechanistically informed design process. Rather than discovering the resulting personality and potential problems through trial and error after deployment, users can proactively identify and mitigate risks during the creation phase.
@@
We conducted a controlled study to evaluate how feedback based on mechanistic interpretability influences users’ comprehension of chatbot behaviors and their prompt refinement strategies. Our study examines:
@@
  • whether neural transparency insight improves user comprehension of their personalized AI, shapes their perception of the model, and helps them achieve their desired system prompt;
@@
  • how accurately users can anticipate their AI’s behavior compared to internal model activations;
@@
    Empirical evidence demonstrating how mechanistic interpretability-based insight improves user trust in chatbots while also being perceived as useful and interactive.
@@ 2.1. Personality and Safety of Personalized AI
As AI systems become increasingly personalized, ensuring safe interaction while preserving user creative control presents a fundamental tension. Some of these current safety techniques include post-training methods that incorporate human and AI feedback on model outputs (christiano2017deep; bai2022constitutional; rafailov2023direct), evaluation against safety benchmarks measuring behaviors such as truthfulness (lin2021truthfulqa) and toxicity (hartvigsen2022toxigen), as well as adversarial testing through red-teaming exercises (ganguli2022red).
However, the non-deterministic nature of LLMs can lead to user-modified prompts producing unpredictable behavior (kim2024persona; fitz2025psychometric). Recent work has shown that training language models to exhibit warm and empathetic personas, traits users commonly desire, undermines the AI’s reliability and increases error rates (ibrahim2025training). The personality given to an LLM using the system prompt has additionally been shown to influence model performance on safety benchmarks (fitz2025psychometric; zheng2024prompt), raising further safety concerns. Current characterization of these chatbot personalities has been done using established frameworks such as the Big Five personality model (jiang2022mpi; jiang2023personallm), but such traits may be less relevant to user preferences. However, existing methods to assess model personality operate at inference time by analyzing model responses, creating two significant limitations: (1) prevention of rapid iteration during system prompt design and (2) requirement of substantial computational resources in each round of inference. Other work has explored user agency with model personality by matching users with pre-defined LLM personas for support (tu2023characterchat) and incorporating role-playing personas to enhance zero-shot reasoning (kim2024persona). Despite these advances, limited research exists on enabling users to design custom personas tailored to their specific needs or on developing interfaces that facilitate this design process.
@@ 2.2. Mechanistic Interpretability
There is increasing evidence that LLMs represent features as linear directions in the representational space created by its activations (elhage2022superposition; nanda2023emergent; zou2023representation). This phenomenon arises from polysemanticity, whereby LLMs encode more features than available neurons, necessitating individual neurons to represent multiple features through linear combinations of their activation values (elhage2022superposition). Various concepts and behaviors, such as refusal (arditi2024refusal), sentiment (tigges2023linear), truth (marks2023geometry), political beliefs (kim2025linear), and spatial and temporal relationships (gurnee2023language) appear to be encoded in this way. Personality traits can also be encoded as linear directions through persona vectors, which can be found by using difference-in-means between contrastive model responses(chen2025persona).
Linear probes (alain2016understanding), which use linear classifiers applied to activations of the model, can similarly identify and manipulate these linearly represented features. Previous work has used linear probes to build human-AI interfaces that measure an LLM’s internal representations of user demographics along these linear feature dimensions (chen2024designing). The interface allowed users to both directly view and manipulate these internal representations and was found to improve transparency and user experience. However, linear probes require extensive data collection and classifier training, whereas persona vectors can be computed from smaller datasets and without requiring model training.
While these methods enable model interpretability, translating their insights into actionable user feedback requires effective visualization. Prior work in using MI techniques to visualize the internal representations and mechanisms of AI includes Neuronpedia (neuronpedia), an open source repository and platform for researchers and others with a technical background. Among the most notable features is one that allows users to explore the features present in the activations of LLMs using sparse autoencoders (SAEs) (cunningham2023sparse). SAEs are also applied in other tools that allow for activation steering (turner2023steering), where the user can directly manipulate the values of the activations to change behavior. Another is circuit tracing (ameisen2025circuit), which allows users to view and analyze the connections between features to reveal the LLM’s internal reasoning process. These tools simplify the application of different methodologies in MI to allow researchers to explore the complex internal representations of LLMs with a user interface. While these tools represent initial attempts to leverage MI for making AI mechanisms more accessible, a critical gap remains: translating these insights for non-technical users to enhance their understanding and control over AI systems.
@@ 3. Methodology
Our neural transparency approach translates insights from mechanistic interpretability into a practical tool for chatbot creation. The core design challenge was making features of the neural activations, typically only analyzed by AI researchers, interpretable and actionable for non-technical users. We also developed additional methodological steps to process the persona vectors so they could be used to calibrate and directly compare the expression of the personality traits. We developed a web-based interface that uses persona vectors extracted from model activations to provide real-time predictions about chatbot behavioral traits before users deploy and interact with their creation.
We chose chatbot companion creation as our application domain as it represents a high-stakes context where behavioral prediction failures can have serious consequences, yet users typically receive no transparency about how their design choices translate into AI behavior. We focused specifically on emotional support chatbots, a use case that balances ecological validity (millions use AI companions for emotional support) with ethical constraints (avoiding explicitly adversarial tasks in a controlled study). This task requires users to balance multiple competing objectives: warmth and empathy without excessive sycophancy, honesty without coldness—making it an ideal testbed for evaluating whether neural transparency helps users navigate complex tradeoffs.
@@ 3.1.1. Generating Prompts for Trait Extraction
The first step to create persona vectors was generating contrastive system prompts and situation questions that extract the selected behavior from our target LLM, Llama-3.2-3B-Instruct (dubey2024llama). We chose eight personality dimensions (empathy, sociality, encouraging, funniness, formality, sycophancy, hallucination, and toxicity), each with opposing trait pairs for a total of 16 traits. The first five were selected for relevance to personality traits a user might desire in an AI chatbot. The final three (sycophancy, hallucination, and toxicity) were selected to inform the user about the safety implications of their system prompt.
We generate five contrastive pairs of system prompts using Claude Haiku 3.5. Each pair contained a positive system prompt that instructs the model to exhibit the selected trait and a negative system prompt that instructs the model to exhibit the opposite of the selected trait. These were used to find the direction in the activation space that control the expression of different levels of the target trait. An example for empathy is the following: